# latexFont shape

### v1

@@ Abstract
Huggingface: https://huggingface.co/TsinghuaC3I
@@ 1 Introduction
Despite the remarkable capabilities, proprietary models may face security and privacy challenges due to the sensitive nature of medical data, such as potential data breaches and the risk of exposing sensitive patient information [35, 79, 40]. On the other hand, open-source LLMs can be customized and adapted to specific healthcare contexts by fine-tuning on local datasets, enabling the development of models tailored to the needs of specific patient populations, healthcare settings, or research questions, thereby enhancing their practical utility and impact. Exploring how to build open-source, GPT-4-level LLMs in the field of biomedicine is underway. Beyond supervised fine-tuning, preference learning technologies like Reinforcement Learning from Human or AI Feedback(RLHF or RLAIF) [33, 29], direct preference optimization (DPO) [52], Kahneman-Tversky Optimization (KTO) [17] and others [73, 23, 24, 43] has proven to play a significant role in enhancing the reasoning abilities of open LLMs in various tasks such as coding, mathematics, and logic [44, 77]. However, preference learning remains under-explored in the biomedical community, which is mainly limited by the scarcity of high-quality and extensively annotated preference datasets.
@@ 2.3 Dataset Statistics
As illustrated in Table 1, the UltraMedical collections ultimately comprise 410K instructions. For the preference annotation, we select the instructions with the highest scores from each dataset, resulting in approximately 100K instructions accompanied by eight models’ completions. During the preference binarization process, we aim to maximize the selection, achieving $C^{2}_{8}=28$ combinations of “chosen” and “rejected” completions per instruction. Although we retain only completions with differing scores, we ultimately obtain approximately 1.8M pairs for reward modeling (approximately 18 times the size of the instruction.). We provide more details in Appendix C.
@@ 6 Conclusion
We express our gratitude to the clinical experts and graduate students majoring in biomedicine from The First Affiliated Hospital of Nanchang University. Their professional knowledge contributed invaluable suggestions towards the construction of UltraMedical, facilitated the correction of annotation errors, and provided a more robust reward benchmark. We also extend our appreciation to the open-source community for sharing dataset sources, which served as essential components of UltraMedical. Furthermore, we acknowledge the contributions of all open-source language model providers, whose efforts have significantly propelled the advancement of research in this domain.
@@ C.2 Details of UltraMedical Preference
We present the model’s accuracy for QA tasks in Figure 7(a), the models’ win percentages in binarized preference in Figure 10, and the scores and rankings of all models across various tasks from GPT-4 in Figures 9 and 8.
@@ C.3 Details of Medical Reward Bench
C.4 Details of General Instructions

### v2

@@ Abstract
Huggingface: https://hf.co/collections/TsinghuaC3I
@@ 1 Introduction
Despite the remarkable capabilities, proprietary models may face security and privacy challenges due to the sensitive nature of medical data, such as potential data breaches and the risk of exposing sensitive patient information [35, 80, 40]. On the other hand, open-source LLMs can be customized and adapted to specific healthcare contexts by fine-tuning on local datasets, enabling the development of models tailored to the needs of specific patient populations, healthcare settings, or research questions, thereby enhancing their practical utility and impact. Exploring how to build open-source, GPT-4-level LLMs in the field of biomedicine is underway. Beyond supervised fine-tuning, preference learning technologies like Reinforcement Learning from Human or AI Feedback(RLHF or RLAIF) [33, 29], direct preference optimization (DPO) [52], Kahneman-Tversky Optimization (KTO) [17] and others [73, 23, 24, 43] has proven to play a significant role in enhancing the reasoning abilities of open LLMs in various tasks such as coding, mathematics, and logic [44, 78]. However, preference learning remains under-explored in the biomedical community [77], which is mainly limited by the scarcity of high-quality and extensively annotated preference datasets.
@@ 2.2.2 Preference Annotation
Human Annotation: To ensure the reliability of the medical reward benchmark, we assembled a team of three experts, each with at least three years of research experience in biomedicine. They utilized a customized WebUI and academic search engines to validate question-answer pairs. For the reward benchmark, out of 1,000 test samples, only about 780 were retained where at least two annotators agreed on the same label. Samples with disagreements or both incorrect answers were removed. We provide more details about human annotation in Appendix C.4.
Annotation Cost: The costs associated with creating the dataset and benchmark primarily include GPT-4-Turbo API (version 1106) calls for instruction synthesis and response generation, as well as preference annotation, totaling approximately \$20,000.
@@ 2.3 Dataset Statistics
Overall: As illustrated in Table 1, the UltraMedical collections ultimately comprise 410K instructions. For the preference annotation, we select the instructions with the highest scores from each dataset, resulting in approximately 100K instructions accompanied by eight models’ completions. During the preference binarization process, we aim to maximize the selection, achieving $C^{2}_{8}=28$ combinations of “chosen” and “rejected” completions per instruction. Although we retain only completions with differing scores, we ultimately obtain approximately 1.8M pairs for reward modeling (approximately 18 times the size of the instruction.). We provide more details in Appendix C.
Medical RewardBench: For the initially given 1,000 test pairs, we ultimately retained 777 pairs following human expert annotation. These include 238 easy, 196 hard, 180 length-based, and 163 human-judged pairs. Approximately 233 pairs were filtered out due to issues such as incorrect formulations, difficulty in answering, or both. The human category comprises pairs where preferences differ between human annotators and GPT-4, which is regraded as even hard for GPT-4 to recognize.
@@ 6 Conclusion
Acknowledgments We express our gratitude to the clinical experts and graduate students majoring in biomedicine from The First Affiliated Hospital of Nanchang University. Their professional knowledge contributed invaluable suggestions towards the construction of UltraMedical, facilitated the correction of annotation errors, and provided a more robust reward benchmark. We also extend our appreciation to the open-source community for sharing dataset sources, which served as essential components of UltraMedical. Furthermore, we acknowledge the contributions of all open-source language model providers, whose efforts have significantly propelled the advancement of research in this domain.
Disclosure of Funding This work is supported by the National Science and Technology Major Project (2023ZD0121403), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), and National Natural Science Foundation of China (No. 62406165).
@@ C.2 Details of UltraMedical Preference
We present the model’s accuracy for QA tasks in Figure 9(a), the models’ win percentages in binarized preference in Figure 12, and the scores and rankings of all models across various tasks from GPT-4 in Figures 11 and 10.
@@ C.3 Details of Medical Reward Bench
C.4 Details of Human Annotation
We requested a human expert to review and re-annotate 1,000 preference pairs. The web interface used for annotation is displayed in Figure 6.
For biomedical-related questions (including clinical, exam, and open research questions) along with responses from Model A and Model B, the task is to choose the best response (vote) based on the following criteria:
  • “Honest”: The response is more accurate and verifiable with no factual errors.
  • “Helpful”: The response is more useful, addresses the problem effectively, and does not contain platitudes.
  • “Harmless”: The response is safe and ethically sound (generally always the case).
  • “Length Bias”: A longer response is not necessarily better; avoid verbosity, prioritizing the above 3H.
Annotation Process:
  • Enter the name of the annotator for later processing and filtering of invalid annotations.
  • Review the question and the answers from the two models (for reference, see GPT-4’s explanation), and vote for the best response.
  • After clicking to vote, the question will automatically refresh. To simplify the process, returning to a previous question is not supported!
  • If a suitable response cannot be selected due to a lack of knowledge, errors in the question or answers, etc., click skip.
  • Also, leave a comment in the comments section with the corresponding question number.
Figure 6: The WebUI used by human experts to annotate and correct preferences.
@@ C.4 Details of Human Annotation
C.5 Details of General Instructions